# MagicDrive3D: Controllable 3D Generation for Any-View Rendering in Street Scenes

### v2

@@ Abstract
While controllable generative models for images and videos have achieved remarkable success, high-quality models for 3D scenes, particularly in unbounded scenarios like autonomous driving, remain underdeveloped due to high data acquisition costs. In this paper, we introduce MagicDrive3D, a novel pipeline for controllable 3D street scene generation that supports multi-condition control, including BEV maps, 3D objects, and text descriptions. Unlike previous methods that reconstruct before training the generative models, MagicDrive3D first trains a video generation model and then reconstructs from the generated data. This innovative approach enables easily controllable generation and static scene acquisition, resulting in high-quality scene reconstruction. To address the minor errors in generated content, we propose deformable Gaussian splatting with monocular depth initialization and appearance modeling to manage exposure discrepancies across viewpoints. Validated on the nuScenes dataset, MagicDrive3D generates diverse, high-quality 3D driving scenes that support any-view rendering and enhance downstream tasks like BEV segmentation. Our results demonstrate the framework’s superior performance, showcasing its transformative potential for autonomous driving simulation and beyond. 00footnotetext: Corresponding authors. Project Page: https://flymin.github.io/magicdrive3d.
Figure 1: Rendered panorama of the street scene generated from MagicDrive3D. With conditional controls from 3D bounding boxes of objects, BEV road map, and text descriptions (e.g., weather), MagicDrive3D generates complex open-world 3D scenes represented by deformable Gaussians.
@@ 1 Introduction
With the advancement of generative models, particularly diffusion models (Goodfellow et al., 2014; Ho et al., 2020; Song et al., 2020; Rombach et al., 2022), there has been increasing interest in generating 3D assets (Poole et al., 2022; Vahdat et al., 2022; Bautista et al., 2022). While a significant amount of work has focused on object-centric generation (Poole et al., 2022; Tang et al., 2024), generating open-ended 3D scenes remains relatively unexplored. This gap is even more critical because many downstream applications, such as Virtual Reality (VR) and autonomous driving simulation, require controllable generation of 3D street scenes, which is an open challenge.
@@
3D-aware view synthesis111 In this paper, we focus on generative models where views/scenes are generated from latent variables. methods can be broadly categorized into two approaches: geometry-free view synthesis and geometry-focused scene generation (Rombach et al., 2021). Geometry-free methods directly generate 2D images (Chen et al., 2024) or videos (Gao et al., 2024; Wen et al., 2023; Wang et al., 2023b) based on camera parameters, excelling in photo-realistic image generation. However, they often lack sufficient geometric consistency, limiting their ability to extend to viewpoints beyond the dataset (Gao et al., 2024; Wen et al., 2023; Wang et al., 2023b). On the other hand, geometry-focused methods (e.g., GAUDI (Bautista et al., 2022) and NF-LDM (Kim et al., 2023)) generate 3D representations (e.g., NeRF (Mildenhall et al., 2020) or voxel grids) from latent inputs, supporting multi-view rendering. Despite their broader applicability, these methods require expensive data collection, necessitating static scenes and consistent sensor properties like exposure and white balance. Street view datasets, such as nuScenes (Caesar et al., 2020), often fail to meet these requirements, making it extremely challenging to train geometry-focused 3D street scene generation models using such datasets.
@@
Recognizing the advancements in controllable generation by geometry-free view synthesis methods (Chen et al., 2024; Gao et al., 2024), it is potential to use them as data engines. Their controllability and photo-realism could address the challenges faced by geometry-focused methods. However, the limited 3D consistency in synthetic views from geometry-free methods, such as temporal inconsistency among frames and deformation of objects, poses crucial issues for integrating two kinds of methods into a unified framework.
@@
To address these challenges, we propose MagicDrive3D, a novel framework that combines geometry-free view synthesis and geometry-focused reconstruction for controllable 3D street scene generation. As illustrated in Figure 2, our approach begins with training a multi-view video generation model to synthesize multiple views of a static scene. This model is configured using controls from object boxes, road maps, text prompts, and camera poses. To enhance inter-frame 3D consistency, we incorporate coordinate embeddings that represent the relative transformation between LiDAR coordinates for accurate control of frame positions. Next, we improve the reconstruction quality of generated views by enhancing 3D Gaussian splatting from the perspectives of prior knowledge, modeling, and loss functions. Given the limited overlap between different camera views (Xie et al., 2023), we adopt a monocular depth prior and propose a dedicated algorithm for alignment in sparse-view settings. Additionally, we introduce deformable Gaussian splatting and appearance embedding maps to handle local dynamics and exposure discrepancies, respectively.
@@
Demonstrated by extensive experiments, our MagicDrive3D framework excels in generating highly realistic street scenes that align with road maps, 3D bounding boxes, and text descriptions, as exemplified in Figure 1. We show that the generated camera views can augment training for Bird’s Eye View (BEV) segmentation tasks, providing comprehensive controls for scene generation and enabling the creation of novel street scenes for autonomous driving simulation. Notably, MagicDrive3D is the first to achieve controllable 3D street scene generation using a common driving dataset (e.g., the nuScenes dataset (Caesar et al., 2020)), without requiring repeated data collection from static scenes.
@@ 2 Related Work
3D Scene Generation. Numerous 3D-aware generative models can synthesize images with explicit camera pose control (Zhao et al., 2024; Rombach et al., 2021) and potentially other scene properties (Tang et al., 2023), but only a few scale for open-ended 3D scene generation. GSN (DeVries et al., 2021) and GAUDI (Bautista et al., 2022), representative of models generating indoor scenes, utilize NeRF (Mildenhall et al., 2020) with latent code input for “floorplan” or tri-plane feature. Their reliance on datasets covering different camera poses is incompatible with typical driving datasets where camera configuration remains constant. NF-LDM (Kim et al., 2023) develops a hierarchical latent diffusion model for scene feature voxel grid generation. However, their representation and complex modeling hinder fine detail generation.
@@
Contrary to previous works focusing on scene generation using explicit geometry, often requiring substantial data not suitable for typical street view datasets (e.g., nuScenes (Caesar et al., 2020)) as discussed in Section 1, we propose merging geometry-free view synthesis with geometry-focused scene representations for controllable street scene creation. Methodologically, LucidDreamer (Chung et al., 2023) is most similar to our approach, although it relies on a text-controlled image generation model, which cannot qualify as a view synthesis model. In contrast, our video generation model is 3D-aware. Besides, we propose several improvements over 3DGS for better scene generation quality. Besides, inpainting-based methods like LucidDreamer (Chung et al., 2023) and WonderJourney (Yu et al., 2024) cannot complete our controllable street scene generation task. We showcase the differences in Appendix C.
@@
Street View Video Generation. Diffusion models (Song et al., 2020; Ho et al., 2020) have influenced a range of works on street view video generation, from single to multi-view videos (e.g., (Wang et al., 2023a; Gao et al., 2024; Wen et al., 2023; Wang et al., 2023b)). Despite cross-view consistency being essential for multi-view video generation, their generalization ability on camera poses is limited due to their data-centric nature (Gao et al., 2024). Furthermore, these models lack exact control over frame transformation (i.e., precise car trajectory), which is crucial for scene reconstruction. Our work addresses this by enhancing control in video generation and proposing a dedicated deformable Gaussian splatting for geometric assurance.
@@
Street Scene Reconstruction. Scene reconstruction and novel view rendering for street views are useful in applications like driving simulation, data generation, and augmented and virtual reality. For street scenes, attributes like scene dynamic and discrepancies from multi-camera data collection make typical large-scale reconstruction methods ineffective (e.g., Rematas et al. (2022); Martin-Brualla et al. (2021); Lin et al. (2024)). Hence, real data-based reconstruction methods like Xie et al. (2023); Yan et al. (2024) utilize LiDAR for depth prior, but their output only permits novel view rendering from the same scene. Unlike these methods, our approach synthesizes novel scenes under multiple levels of conditional controls.
@@ 3 Preliminaries
Problem Formulation. In this paper, we focus on controllable street scene generation. Given scene description $\mathbf{S}_{t}$, our task is to generate street scenes (represented with 3D Gaussians $\mathbf{G}$) that correspond to the description from a set of latent $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, i.e. $\mathbf{G}=\mathcal{G}(\mathbf{S}_{t},\mathbf{z})$. To describe a street scene, we adopt the most commonly used conditions as per Gao et al. (2024); Wang et al. (2023b); Wen et al. (2023). Specifically, a frame of driving scene $\mathbf{S}_{t}=\{\mathbf{M}_{t},\mathbf{B}_{t},\mathbf{L}_{t}\}$ is described by road map $\mathbf{M}_{t}\in\{0,1\}^{w\times h\times c}$ (a binary map representing a $w\times h$ meter road area in BEV with $c$ semantic classes), 3D bounding boxes $\mathbf{B}_{t}=\{(c_{i},b_{i})\}_{i=1}^{N}$ (each object is described by box $b_{i}=\{(x_{j},y_{j},z_{j})\}_{j=1}^{8}\in\mathbb{R}^{8\times 3}$ and class $c_{i}\in\mathcal{C}$), and text $\mathbf{L}_{t}$ describing additional information about the scene (e.g., weather and time of day). In this paper, we parameterize all geometric information according to the LiDAR coordinate of the ego car.
One direct application of scene generation is any-view rendering. Specifically, given any camera pose $\mathbf{P}=[\mathbf{K},\mathbf{R},\mathbf{t}]$ (i.e., intrinsics, rotation, and translation), the model $\mathbf{G}(\cdot)$ should render photo-realistic views with 3D consistency, $\mathcal{I}^{r}=\mathbf{G}(\mathbf{P})$, which is not applicable to previous street view generation (e.g., Gao et al. (2024); Wang et al. (2023b); Wen et al. (2023)). Besides, we present more applications in Section 5.
3D Gaussian Splatting. We briefly introduce 3DGS since our scene representation is based on it. 3DGS (Kerbl et al., 2023) represents the geometry and appearance via a set of 3D Gaussians $\mathbf{G}$. Each 3D Gaussian is characterized by its position $\boldsymbol{\mu}_{p}$, anisotropic covariance $\boldsymbol{\Sigma}_{p}$, opacity $\alpha_{p}$, and spherical harmonic coefficients for view-dependent colors $c_{p}$. Given a sparse point cloud $\mathcal{P}$ and several camera views $\{\mathcal{I}_{i}\}$ with poses $\{\mathbf{P}_{i}\}$, a point-based volume rendering (Zwicker et al., 2001) is applied to make Gaussians optimizable through gradient descend and interleaved point densification. Specifically, the loss is as follows:
$\mathcal{L}_{\text{GS}}=(1-\lambda)\mathcal{L}_{1}(\mathcal{I}^{r}_{i},\mathcal{I}_{i})+\lambda\mathcal{L}_{\text{D-SSIM}}(\mathcal{I}^{r}_{i},\mathcal{I}_{i}),$ (1)
where $\mathcal{I}^{r}$ is the rendered image, $\lambda$ is a hyper-parameter, and $\mathcal{L}_{\text{D-SSIM}}$ denotes the D-SSIM loss (Kerbl et al., 2023).
@@ 4 Methods
Direct modeling of controllable street scene generation faces two major challenges: scene dynamics and discrepancy in data collection. Scene dynamics refer to the movements and deformation of elements in the scene, while discrepancy in data collection refer to the discrepancy (e.g., exposure) caused by data collection. These two challenges are even more severe due to the sparsity of cameras for street views (e.g., typically only 6 surrounding perspective cameras). Therefore, reconstruction-generation frameworks do not work well for street scene generation (Kim et al., 2023; Bautista et al., 2022).
@@
MagicDrive3D integrate geometry-free view synthesis and geometry-focused reconstruction, where control signals are tackled by a multi-view video generator, while reconstruction step guarantee the generalization ability for any-view rendering. Such a video generator has two advantages: first, since multi-view video generation does not require generalization on novel views (Gao et al., 2024), it poses less data dependency for street scenes; second, through conditional training, the model is capable of decomposition of control signals, and thus turns dynamic scenes into static scenes which are more friendly for reconstruction. Besides, for the reconstruction step, strong prior from the multi-view video reduces the burden for scene modeling with complex details.
@@ 4.2 Relative Pose Control for Video Generation
Given scene descriptions and a sequence of camera poses $\{\mathbf{S}_{t},\mathbf{P}_{c,t}\}$, our video generator is responsible for multi-view video generation. Although many previous art for street view generation achieve expressive visual effects, such as Gao et al. (2024); Wen et al. (2023); Wang et al. (2023b; a), their formulations leave out a crucial requirement for 3D modeling. Specifically, the camera pose $\mathbf{P}_{c,t}$ is typically relative to the LiDAR coordinate of each frame. Thus, there is no precise control signal related to the ego trajectory, which significantly determine the geometric relationship between views of different $t$s.
In our video generation model, we amend such precise control ability by adding the transformation between each frame to the first frame, i.e., $\mathbf{T}^{0}_{t}$. To properly encode such information, we adopt Fourier embedding with Multi-Layer Perception (MLP), and concatenate the embedding with the original embedding of $\mathbf{P}_{c,t}$, similar to Gao et al. (2024). As a result, our video generator provides better 3D consistency across frames, most importantly, making the camera poses to each view available in the same coordinate, i.e., $[\mathbf{R}^{0}_{c,t},\mathbf{t}^{0}_{c,t}]=\mathbf{T}^{0}_{t}[\mathbf{R}_{c,t},\mathbf{t}_{c,t}]$.
4.3 Enhanced Gaussian Splatting for Generated Content
@@
As introduced in Section 3, 3DGS is a flexible explicit representation for scene reconstruction. Besides, the fast training and rendering speed of 3DGS make it highly suitable for reducing generation costs in our scene creation pipeline. However, similar to other 3D reconstruction methods, 3DGS necessitates high cross-view 3D consistency at the pixel level, which unavoidably magnifies the minute errors in the generated data into conspicuous artifacts. Therefore, we propose improvements for 3DGS from the perspectives of prior, modeling, and loss, enabling 3DGS to tolerate minor errors in the generated camera view, thereby becoming a potent tool for enhancing geometric consistency in rendering.
Prior: Consistent Depth Prior. As essential geometry information, depth is extensively utilized in street scene reconstruction, such as the depth value from LiDAR or other depth sensors used in Xie et al. (2023); Yan et al. (2024). However, for synthesized camera views, the depth is unavailable. Therefore, we propose to use pre-trained monocular depth estimator (Bhat et al., 2023) to infer depth information.
While monocular depth estimation is separate for each camera view, proper scale $s_{c,t}$ and offset $b_{c,t}$ parameters should be estimated to align them for a single scene (Zhou et al., 2023), as in Figure 3(a). To this end, we first apply the Point Cloud (PCD) from Structure of Motion (SfM) (Schönberger et al., 2016; Schönberger & Frahm, 2016), shown in Figure 3(b). However, such PCD is too sparse to accurately restore $(s_{c,t},b_{c,t})$ for any views. To bridge the final gap, secondly, we propose further optimizing the $(s_{c,t},b_{c,t})$ using the GS loss, as in Figure 3(c). Specifically, we replace the optimization for Guassian centers $\boldsymbol{\mu}_{i}$ with $(s_{c,t},b_{c,t})$. After the optimization, we initialize $\boldsymbol{\mu}_{i}$ with points from depth values. Since GS algorithm is sensitive to accurate point initialization (Kerbl et al., 2023; Fan et al., 2024), our method provides useful prior to reconstructing in this sparse view scenario.
Modeling: Deformable Gaussians for Local Dynamic. Despite the 3D geometric consistency provided by our video generation model, there are inevitably pixel-level disparities in some object details, as shown in Figure 4. The strict consistency assumption of 3DGS may amplify these minor errors, resulting in floater artifacts. To mitigate the impact of these errors, we propose Deformable Gaussian Splitting (DGS), which, based on 3DGS, reduces the requirement for temporal consistency between frames, thereby ensuring the reconstruction effect of the generated viewpoint.

### v3

@@ Abstract
While controllable generative models for images and videos have achieved remarkable success, high-quality models for 3D scenes, particularly in unbounded scenarios like autonomous driving, remain underdeveloped due to high data acquisition costs. In this paper, we introduce MagicDrive3D, a novel pipeline for controllable 3D street scene generation that supports multi-condition control, including BEV maps, 3D objects, and text descriptions. Unlike previous methods that reconstruct before training the generative models, MagicDrive3D first trains a video generation model and then reconstructs from the generated data. This innovative approach enables easily controllable generation and static scene acquisition, resulting in high-quality scene reconstruction. To address the minor errors in generated content, we propose deformable Gaussian splatting with monocular depth initialization and appearance modeling to manage exposure discrepancies across viewpoints. Validated on the nuScenes dataset, MagicDrive3D generates diverse, high-quality 3D driving scenes that support any-view rendering and enhance downstream tasks like BEV segmentation. Our results demonstrate the framework’s superior performance, showcasing its potential for autonomous driving simulation and beyond.
Figure 1: Rendered panorama of the street scene generated from MagicDrive3D. With conditional controls from 3D bounding boxes of objects, BEV road map, ego trajectory, and text descriptions (e.g., timeofday), MagicDrive3D generates complex open-world 3D scenes represented by deformable Gaussians.
00footnotetext: Corresponding authors.
Project Page: https://flymin.github.io/magicdrive3d/.
@@ 1 Introduction
With the advancement of generative models [10, 11, 25, 22], there has been increasing interest in generating 3D assets [19, 28, 2]. While many works focus on object-centric generation [19, 27], generating open-ended 3D scenes remains relatively unexplored. This gap is even more critical considering many downstream applications, such as virtual reality and autonomous driving, require controllable generation of 3D street scenes, which is an open challenge.
@@
The typical workflow for 3D scene generation is geometry-focused [^1] In this paper, we focus on generative models where views/scenes are generated from latent variables.. It begins with reconstructing the scene from multi-view sensors to obtain a compact representation (e.g., voxel grids in NF-LDM [13] and tri-planes in GAUDI [2]), followed by training generative models to produce the scene representation from latent variables. Despite their broad applicability, these methods require expensive data collection for the initial reconstruction step, which sets the upper limit on generation quality. Specifically, they often require static scenes and consistent sensor properties, such as exposure and white balance. Street view datasets, like nuScenes [4], fail to meet these requirements, making it extremely challenging to train geometry-focused 3D street scene generation models using such datasets.
Geometry-free methods [21] represent another category in 3D-aware view synthesis. In street-view generation, multi-view 2D images or videos can be directly generated based on camera parameters [9, 32, 30], excelling in photorealistic image generation. However, they often lack sufficient geometric consistency, limiting their ability to extend to viewpoints beyond the dataset [9, 32, 30].
@@
Recognizing the advancements in controllable generation by geometry-free view synthesis methods [5, 9], it is potential to use them as data engines. Their controllability and photo-realism could address the challenges faced by geometry-focused methods. However, the limited 3D consistency in synthetic views from geometry-free methods, such as temporal inconsistency among frames and deformation of objects, poses crucial issues for integrating two kinds of methods into a unified framework.
@@
To address these challenges, we propose MagicDrive3D, a novel framework that combines geometry-free view synthesis and geometry-focused reconstruction for controllable 3D street scene generation. As illustrated in Figure 2, our approach begins with training a multi-view video generation model to synthesize multiple views of a static scene. This model is configured using controls from object boxes, road maps, text prompts, and camera poses. To enhance inter-frame 3D consistency, we incorporate coordinate embeddings that represent the relative transformation between LiDAR coordinates for accurate control of frame positions. Next, we improve the reconstruction quality of generated views by enhancing 3D Gaussian splatting from the perspectives of prior knowledge, modeling, and loss functions. Given the limited overlap between different camera views [34], we adopt a monocular depth prior and propose a dedicated algorithm for alignment in sparse-view settings. Additionally, we introduce deformable Gaussian splatting and appearance embedding maps to handle local dynamics and exposure discrepancies, respectively.
@@
Demonstrated by extensive experiments, our MagicDrive3D framework excels in generating highly realistic street scenes that align with road maps, 3D bounding boxes, ego trajectory, and text descriptions, as exemplified in Figure 1. We show that the generated camera views can augment training for Bird’s Eye View (BEV) segmentation tasks, improving the view-point robustness on perception tasks. Notably, MagicDrive3D is the first to achieve controllable 3D street scene generation using a common driving dataset (e.g., the nuScenes dataset [4]), without requiring repeated data collection from static scenes.
@@ 2 Related Work
3D Scene Generation. Numerous 3D-aware generative models can synthesize images with explicit camera pose control [38, 21] and potentially other scene properties [26], but only a few scale for open-ended 3D scene generation. GSN [7] and GAUDI [2], representative of models generating indoor scenes, utilize NeRF [18] with latent code input for “floorplan” or tri-plane feature. Their reliance on datasets covering different camera poses is incompatible with typical driving datasets where camera configuration remains constant. NF-LDM [13] develops a hierarchical latent diffusion model for scene feature voxel grid generation. However, their representation and complex modeling hinder fine detail generation.
@@
Contrary to previous works focusing on scene generation using explicit geometry, often requiring substantial data not suitable for typical street view datasets (e.g., nuScenes [4]) as discussed in Section 1, we propose merging geometry-free view synthesis with geometry-focused scene representations for controllable street scene creation. Methodologically, LucidDreamer [6] is most similar to our approach, although it relies on a text-controlled image generation model, which cannot qualify as a view synthesis model. In contrast, our video generation model is 3D-aware. Besides, we propose several improvements over 3DGS for better scene generation quality. Besides, inpainting-based methods like LucidDreamer [6] and WonderJourney [36] cannot complete our controllable street scene generation task. We showcase the differences in Appendix E.
@@
Street View Video Generation. Diffusion models [25, 11] boost a series of works for street view video generation. These works span from single view video (e.g., [29]) to multi-view video (e.g., [9, 32, 30]). Despite cross-view consistency being essential for multi-view video generation, their generalization ability on camera poses is limited due to the data-centric nature [9]. Furthermore, these models lack exact control over frame transformation (i.e., precise ego trajectory), which is crucial for scene reconstruction. Our work addresses this by enhancing control in video generation and proposing a dedicated deformable Gaussian splatting for geometric assurance.
@@
Street Scene Reconstruction. Scene reconstruction and novel view rendering for street views are useful in applications like driving simulation, data generation, and augmented and virtual reality. For street scenes, attributes like scene dynamic and discrepancies from multi-camera data collection make typical large-scale reconstruction methods ineffective (e.g., [20, 16, 15]). Hence, real data-based reconstruction methods like [34, 35] utilize LiDAR for depth prior, but their output only permits novel view rendering from the same scene. Unlike these methods, our approach synthesizes novel scenes under multiple levels of conditional controls.
@@ 3 Preliminaries
Problem Formulation. In this paper, we focus on controllable street scene generation. Given scene description $\mathbf{S}_{t}$, our task is to generate street scenes (represented with 3D Gaussians $\mathbf{G}$) that correspond to the description from a set of latent $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, i.e. $\mathbf{G}=\mathcal{G}(\mathbf{S}_{t},\mathbf{z})$. To describe a street scene, we adopt the most commonly used conditions as per [9, 30, 32]. Specifically, a frame of driving scene $\mathbf{S}_{t}=\{\mathbf{M}_{t},\mathbf{B}_{t},\mathbf{L}_{t}\}$ is described by road map $\mathbf{M}_{t}\in\{0,1\}^{w\times h\times c}$ (a binary map representing a $w\times h$ meter road area in BEV with $c$ semantic classes), 3D bounding boxes $\mathbf{B}_{t}=\{(c_{i},b_{i})\}_{i=1}^{N}$ (each object is described by box $b_{i}=\{(x_{j},y_{j},z_{j})\}_{j=1}^{8}\in\mathbb{R}^{8\times 3}$ and class $c_{i}\in\mathcal{C}$), and text $\mathbf{L}_{t}$ describing additional information about the scene (e.g., weather and time of day). In this paper, we parameterize all geometric information according to the LiDAR coordinate of the ego car.
One direct application of scene generation is any-view rendering. Specifically, given any camera pose $\mathbf{P}=[\mathbf{K},\mathbf{R},\mathbf{t}]$ (i.e., intrinsics, rotation, and translation), the model $\mathbf{G}(\cdot)$ should render photo-realistic views with 3D consistency, $\mathcal{I}^{r}=\mathbf{G}(\mathbf{P})$, which is not applicable to previous street view generation (e.g., [9, 30, 32]). Besides, we present more applications in Section 5.
3D Gaussian Splatting. We briefly introduce 3DGS since our scene representation is based on it. 3DGS [12] represents the geometry and appearance via a set of 3D Gaussians $\mathbf{G}$. Each 3D Gaussian is characterized by its position $\boldsymbol{\mu}_{p}$, anisotropic covariance $\boldsymbol{\Sigma}_{p}$, opacity $\alpha_{p}$, and spherical harmonic coefficients for view-dependent colors $c_{p}$. Given a sparse point cloud $\mathcal{P}$ and several camera views $\{\mathcal{I}_{i}\}$ with poses $\{\mathbf{P}_{i}\}$, a point-based volume rendering [41] is applied to make Gaussians optimizable through gradient descend and interleaved point densification. Specifically, the loss is as Equation 1, where $\mathcal{I}^{r}$ is the rendered image, $\lambda$ is a hyper-parameter, and $\mathcal{L}_{\text{D-SSIM}}$ denotes the D-SSIM loss [12].
$\mathcal{L}_{\text{GS}}=(1-\lambda)\mathcal{L}_{1}(\mathcal{I}^{r}_{i},\mathcal{I}_{i})+\lambda\mathcal{L}_{\text{D-SSIM}}(\mathcal{I}^{r}_{i},\mathcal{I}_{i})$ (1)
@@ 4 Methods
Direct modeling of controllable street scene generation faces two major challenges: scene dynamics and discrepancy in data collection. Scene dynamics refer to the movements and deformation of elements in the scene, while discrepancy in data collection refer to the discrepancy (e.g., exposure) caused by data collection. These two challenges are even more severe due to the sparsity of cameras for street views (e.g., typically only 6 surrounding perspective cameras). Therefore, reconstruction-generation frameworks do not work well for street scene generation [13, 2].
@@
MagicDrive3D integrates geometry-free view synthesis and geometry-focused reconstruction, where control signals are tackled by a multi-view video generator, while the reconstruction step guarantees the generalization ability for any-view rendering. Such a video generator has two advantages: first, since multi-view video generation does not require generalization on novel views [9], it poses less data dependency for street scenes; Second, through conditional training, the model is able to decompose control signals, turning dynamic scenes into static scenes that are more friendly for reconstruction. Besides, for the reconstruction step, strong prior from the multi-view video reduces the burden for scene modeling with complex details.
@@ 4.2 Relative Pose Control for Video Generation
Given scene descriptions and a sequence of camera poses $\{\mathbf{S}_{t},\mathbf{P}_{c,t}\}$, our video generator is responsible for multi-view video generation. Although many previous art for street view generation achieve expressive visual effects, such as [9, 32, 30, 29], their formulations leave out a crucial requirement for 3D modeling. Specifically, the camera pose $\mathbf{P}_{c,t}$ is typically relative to the LiDAR coordinate of each frame. Thus, there is no precise control signal related to the ego trajectory, which significantly determines the geometric relationship between views of different $t$s.
In our video generation model, we amend such precise control ability by adding the transformation between each frame to the first frame, i.e., $\mathbf{T}^{0}_{t}$. To properly encode such information, we adopt Fourier embedding with Multi-Layer Perception (MLP), and concatenate the embedding with the original embedding of $\mathbf{P}_{c,t}$, similar to [9]. As a result, our video generator provides better 3D consistency across frames, most importantly, making the camera poses to each view available in the same coordinate, i.e., $[\mathbf{R}^{0}_{c,t},\mathbf{t}^{0}_{c,t}]=\mathbf{T}^{0}_{t}[\mathbf{R}_{c,t},\mathbf{t}_{c,t}]$.
@@
4.3 Enhanced GS for Generated Content
As introduced in Section 3, 3DGS is a flexible explicit representation for 3D reconstruction. Besides, the fast training and rendering speed of 3DGS make it highly suitable for reducing generation costs in our scene creation pipeline. However, similar to other 3D reconstruction methods, 3DGS necessitates high cross-view 3D consistency at the pixel level, which unavoidably magnifies the minute errors in the generated data into conspicuous artifacts. Therefore, we propose improvements for 3DGS from the perspectives of prior, modeling, and loss, enabling 3DGS to tolerate minor artifacts in the generated camera view from our video generation model, thereby becoming a potent tool for enhancing geometric consistency in rendering.
Prior: Consistent Depth Prior. As essential geometry information, depth is extensively utilized in street scene reconstruction, such as the depth value from LiDAR or other depth sensors used in [34, 35]. However, for synthesized camera views, LiDAR/depth information is unavailable (also making these methods not suitable for our scenario). Therefore, we propose to use pre-trained monocular depth estimator [3] to infer depth information.
While monocular depth estimation is separate for each camera view, proper scale $s_{c,t}$ and offset $b_{c,t}$ parameters should be estimated to align them for a single scene [40], as in Figure 3(a). To this end, we first apply the Point Cloud (PCD) from Structure of Motion (SfM) [24, 23], shown in Figure 3(b). However, such PCD is too sparse to accurately restore $(s_{c,t},b_{c,t})$ for any views. To bridge the final gap, secondly, we propose further optimizing the $(s_{c,t},b_{c,t})$ using the GS loss, as in Figure 3(c). Specifically, we replace the optimization for Guassian centers $\boldsymbol{\mu}_{i}$ with $(s_{c,t},b_{c,t})$. After the optimization, we initialize $\boldsymbol{\mu}_{i}$ with points from depth values. Since GS algorithm is sensitive to accurate point initialization [12, 8], our method provides useful prior to reconstructing in this sparse view scenario.
Modeling: Deformable Gaussians for Local Dynamic. Despite the 3D geometric consistency provided by our video generation model, there are inevitably pixel-level disparities in some object details, as shown in Figure 4. The strict consistency assumption of 3DGS may amplify these minor errors, resulting in floater artifacts. To mitigate the impact of these errors, we propose Deformable Gaussian Splitting (DGS), which, reduces the requirement for temporal consistency between frames, thereby ensuring the reconstruction quality of the generated views.